OpenStudy (anonymous):

2log(x+3)+3log(x-7)-5log(x-2)-log(x^2) what is another way i can write this --- i have an idea but i dont know if i am right or wrong

OpenStudy (anonymous):

\[ \frac{ (x+3)^{2} (x-7)^{3} }{ (x-2)^{5} x ^{^{2}} }\]

OpenStudy (anonymous):

wow that took forever to write that is what i would assume would be it

OpenStudy (amistre64):

\[y=2log(x+3)+3log(x-7)-5log(x-2)-log(x^2)\] \[e^y=e^{2log(x+3)+3log(x-7)-5log(x-2)-log(x^2)}\] \[\large e^y=e^{2log(x+3)}e^{3log(x-7)}e^{-5log(x-2)}e^{-log(x^2)}\] \[\large e^y=\frac{(x+3)^2(x-7)^3}{(x-2)^5(x^2)}\] \[y=log\left[\frac{(x+3)^2(x-7)^3}{(x-2)^5(x^2)}\right]\]

OpenStudy (amistre64):

you did fine, but you missed the log

OpenStudy (anonymous):

ohhhhh ok n_n yay close ... i just need one thing soo close hahah thank you ! @amistre64

OpenStudy (amistre64):

youre welcome